# Near-Optimal Channel Estimation for Dense Array Systems

### v1

@@ VI-B Performance Evaluation Under Perfect Kernel
In this subsection, we use the perfect kernel, ${\bm{\Sigma}}_{{\mathbf{h}}}$, to trigger the water-filling, ice-filling, and MM algorithms, and then perform the Bayesian regression. The NMSE performance as a function of SNR is evaluated in Fig. 5. The SNR ranges from $-10{\>}{\text{dB}}\sim 15{\>}{\text{dB}}$. Thanks to the careful design of observation matrix, the optimal water-filling, the proposed ice-filling, and the proposed MM algorithms remarkably outperform existing benchmarks. In particular, a 5 dB NMSE gap between ice-filling and MMSE, as well as a 15 dB gap between ice-filling and AMP, are visible when the SNR is 10 dB. Notably, the NMSE performance attained by the ice-filling method tightly aligns with that of the ideal water-filling scheme, given that the ice-filling intrinsically allocates $Q$ quantized units of power to the $Q$-timeslot pilots. Another finding of interest is that the NMSE gap between the MM and ice-filling algorithms is no higher than 1 dB, demonstrating the near-optimality of the proposed MM algorithm in phase-only-controllable design.
@@ VI-C Performance Evaluation Under Imperfect Kernel
In some scenarios where the perfect prior kernel $\bm{\Sigma}_{\bf h}$ cannot be obtained, the statistical kernel and artificial kernels defined in Section IV-E can be utilized to replace the perfect kernel $\bm{\Sigma}_{\bf h}$ in Algorithms 1$\sim$3. We first consider the statistical kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{{\mathbf{h}}}{\mathbf{I}}_{M}$. The NMSE as a function of the kernel estimation error $\sigma^{2}_{{\mathbf{h}}}$ is shown in Fig. 8. One can observe that all kernel-dependent methods, including the MMSE, Random ${\mathbf{W}}$, and the proposed estimators, suffer from visible performance loss due to the imperfect kernel input. Fortunately, the proposed methods still hold the superiority among all schemes. In particular, the estimation accuracy of the proposed methods is almost not influenced when $\sigma^{2}_{{\mathbf{h}}}<-15\,{\text{dB}}$, which shows their high robustness to the kernel error. One more interesting finding is that, when $\sigma^{2}_{{\mathbf{h}}}$ increases, the gap between MMSE and our proposed methods, become small. The reason is that, an increasing $\sigma^{2}_{{\mathbf{h}}}$ gradually obscures the structural characteristics of ${\bf{\Sigma}}_{{\bf h}}$, which makes the input kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$ behaves more like an identity matrix. In this case, the importance of designing the observation matrix $\bf W$ is reduced.
@@
We further evaluate the achievable NMSE performance under the artificial kernels $\bm{\Sigma}_{\rm exp}$ and ${\bm{\Sigma}}_{\rm Bes}$. Their hyper-parameters, $\eta_{1}$ and $\eta_{2}$, are determined by the binary search through several numerical experiments, which are set to $\eta_{1}=0.56$ and $\eta_{2}=0.85$ in the considered scenario. In this way, we obtain the NMSE versus the SNR and that versus the pilot length $Q$ in Fig. 9 and Fig. 10, respectively. From these two figures one find that, for a given kernel, the curves of “optimal water-filling”, “proposed ice-filling”, and “proposed MM” are very close, which indicates that these three schemes have similar robustness to imperfect kernels. This phenomenon proves the feasibility of using experiential kernels.
@@ VII Conclusions
Appendix A Derivation of $I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({{\mathbf{y}}_{t};{\mathbf{h}}})$
The MI $I({\mathbf{y}}_{t+1};{\mathbf{h}})$ can be decoupled as
$I({\mathbf{y}}_{t+1};{\mathbf{h}})$
$=\log_{2}\left|{\mathbf{I}}_{t+1}+\frac{1}{\sigma^{2}}{\mathbf{W}}_{t+1}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t+1}\right|$
$=\log_{2}\left|\begin{array}[]{cc}{\mathbf{I}}_{t}+\frac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t}&\frac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{w}}_{t+1}\\ \frac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t}&1+\frac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{w}}_{t+1}\end{array}\right|$ (49)
=$\overset{(a)}{=}\log_{2}\left|\begin{array}[]{cc}{\mathbf{I}}_{t}+\frac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t}&\frac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{w}}_{t+1}\\ {\mathbf{0}}&1+\frac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{t}{\mathbf{w}}_{t+1}\end{array}\right|$ (52)
$=I({\mathbf{y}}_{t};{\mathbf{h}})+\log_{2}\left(1+\frac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{t}{\mathbf{w}}_{t+1}\right),$ (53)
where (a) holds because of the relationship in (56). Clearly, $I({\mathbf{y}}_{t+1};{\mathbf{h}})-I({{\mathbf{y}}_{t};{\mathbf{h}}})$ is $\log_{2}\left(1+\frac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{t}{\mathbf{w}}_{t+1}\right)$.
$\left(\begin{array}[]{cc}{\mathbf{I}}_{t}+\dfrac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t}&\dfrac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{w}}_{t+1}\\ {\mathbf{0}}&1+\dfrac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{t}{\mathbf{w}}_{t+1}\end{array}\right)=$ (56)
$\quad\quad\quad\quad\quad\quad\left(\begin{array}[]{cc}{\mathbf{I}}_{t}&{\mathbf{0}}\\ -\dfrac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t}({\mathbf{I}}_{t}+\dfrac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t})^{-1}&1\end{array}\right)\left(\begin{array}[]{cc}{\mathbf{I}}_{t}+\dfrac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t}&\dfrac{1}{\sigma^{2}}{\mathbf{W}}_{t}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{w}}_{t+1}\\ \dfrac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{W}}_{t}&1+\dfrac{1}{\sigma^{2}}{\mathbf{w}}_{t+1}^{H}{\bm{\Sigma}}_{{\mathbf{h}}}{\mathbf{w}}_{t+1}\end{array}\right)$ (61)
Appendix B Proof of Lemma 1
@@
where $(a)$ comes from the definition of ${\bf\Sigma}_{t-1}$, i.e., ${\bf\Sigma}_{t-1}={{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}$. Then, by substituting (72) and (B) into (64), the inverse of ${\bf{K}}_{t}$ can be written as (76).
${\bf{K}}_{t}^{-1}=\left[\begin{array}[]{cc}{\bf{K}}_{t-1}^{-1}+\dfrac{{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}&-\dfrac{{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}\\ -\dfrac{{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}&\dfrac{1}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}\end{array}\right]$ (76)
@@
It is straightforward to verify that the sum of ${\bf E}$, ${\bf F}$, ${\bf G}$, and ${\bf H}$ is exactly $-{{\bf{\Sigma}}_{t-1}}{\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{t-1}}$. By substituting ${\bf E}$, ${\bf F}$, ${\bf G}$, and ${\bf H}$ into (B), we have
@@
Appendix C Proof of Theorem 2
@@
where the inequality (a) holds because $\left(\beta-\frac{\sigma^{2}}{\lambda_{k}}\right)^{+}\geq\beta-\frac{\sigma^{2}}{\lambda_{k}}$. Combining equations (79) and (C), we arrive at
@@
We introduce the following Lemma 9 to evaluate the inequality (81).
Lemma 9
For the $k$-th eigenvalue $\lambda_{k}^{t}$ obtained at the $t$-th timeslot, if $n_{k}^{t}>0$, we have the inequality
@@
Comparing (81) and (82), it is clear that (81) contradicts Lemma 9 given that $n_{k}\geq p_{k}+1>0$. Therefore, $n_{k}$ should be smaller than $p_{k}+1$.
@@
which contradicts Lemma 9 because $n_{k^{\prime}}>p_{k^{\prime}}\geq 0$. As a result, $n_{k}$ should be larger than $p_{k}-1$, and thereby we get
@@
Appendix D Proof of Lemma 3
@@
Appendix E Proof of Lemma 4
@@
Appendix F Proof of Theorem 3
@@
Appendix G Proof of Lemma 5
@@
Therefore, by substituting (G) into (G), the MSE $\delta_{\rm rnd}$ is given by
@@
Appendix H Proof of Lemma 6
@@
Appendix I Proof of Corollary 1
Observing (36), one can find that the MSE ${\hat{\delta}}$ is a quadratic form w.r.t matrix $\bm{\Pi}$. Thus, the lower-bound of the MSE can be achieved by finding a $\bm{\Sigma}$ such that $\frac{{\partial{\hat{\delta}}}}{{\partial{\bf{\Pi}}}}={\bf 0}_{Q\times M}$ holds. According to (36), the partial derivative of ${\hat{\delta}}$ w.r.t $\bm{\Pi}$ can be derived as
$\frac{{\partial{\hat{\delta}}}}{{\partial{\bf{\Pi}}}}={\left({{{\bf{W}}^{H}}{\bf{{\bf{\Sigma}}_{\bf{h}}W}}+{\sigma^{2}}{\bf{I}}_{Q}}\right)^{*}}{{\bf{\Pi}}^{*}}-{\left({{\bf{W}}^{H}}{{{{\bf{\Sigma}}_{\bf{h}}}}}\right)^{*}}.$ (102)
By letting $\frac{{\partial{\hat{\delta}}}}{{\partial{\bf{\Pi}}}}={\bf 0}_{PM\times N}$ and substituting ${\bf{\Pi}}={\left({{{\bf{W}}^{H}}{\bf{\Sigma W}}+{\sigma^{2}}{\bf{I}}_{Q}}\right)^{-1}}{{\bf{W}}^{H}}{\bf{\Sigma}}$ into (102), the original proof can be reduced to prove that
${\left({{{\bf{W}}^{H}}{{{\bf{\Sigma}}_{\bf{h}}{\bf W}}}+{\sigma^{2}}{\bf{I}}_{Q}}\right)^{-1}}$ ${{\bf{W}}^{H}}{{{\bf{\Sigma}}_{\bf{h}}}}=$
${\left({{{\bf{W}}^{H}}{\bf{\Sigma W}}+{\sigma^{2}}{\bf{I}}_{Q}}\right)^{-1}}{{\bf{W}}^{H}}{\bf{\Sigma}}.$ (103)
Since the proposed method does not depend on the value of noise power, (I) should hold for arbitrary $\sigma^{2}$. By considering the special case when $\sigma^{2}\to\infty$, (I) is equivalent to
${{\bf{W}}^{H}}\left({{{{\bf{\Sigma}}_{\bf{h}}}}-{\bf{\Sigma}}}\right)={{\bf{0}}_{Q\times M}}$ (104)
Furthermore, since our derivation also does not impose the specific form of ${\bf{W}}$, ${\bf{\Sigma}}={{\bf{\Sigma}}_{\bf{h}}}$ is necessary to make the equation always hold. By substituting ${\bf{\Sigma}}={{\bf{\Sigma}}_{\bf{h}}}$ into (36), the lower-bound of the achievable MSE is exactly the trace of the posterior kernel ${{\bf{\Sigma}}}_{{\mathbf{h}}|{\mathbf{y}}}$, rewritten as
$\mathop{\min}\limits_{{\bm{\mu}},{\bm{\Sigma}}}$ $~{}{\hat{\delta}}={\mathsf{Tr}}\left({{{\bf{\Sigma}}_{\bf{h}}}}\right)-$
${\mathsf{Tr}}\left({{{\bf{\Sigma}}_{\bf{h}}}{\bf{W}}{{\left({{{\bf{W}}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{W}}+{\sigma^{2}}{\bf{I}}_{Q}}\right)}^{-1}}{{\bf{W}}^{H}}{{\bf{\Sigma}}_{\bf{h}}}}\right),$ (105)
which completes the proof.
Appendix J Proof of Lemma 7
@@ Appendix I Proof of Corollary 1
which is equal to ${\bm{\Omega}}{{\bf{\hat{\Sigma}}}}_{\bf{h}}$ and obviously the is a function of ${\bf W}{\bf W}^{H}$. Using the same matrix techniques as (J), we can prove that
@@
Appendix K Proof of Lemma 8
Let ${\bf\Sigma_{h}}={\mathbf{U}}{\bf\Lambda}{\mathbf{U}}^{H}$ denote the complete eigenvalue decomposition of the kernel ${{\bf{\Sigma}}_{\bf{h}}}$, wherein ${\mathbf{U}}\in{\mathbb{C}}^{M\times M}$ and ${\mathbf{\Lambda}}=\text{diag}\{\lambda_{1},\cdots,\lambda_{M}\}$ with $\lambda_{m}=0$ for $K<m\leq M$. According to the analyses in Appendices D, E, and G, the term ${\bf W}{\bf W}^{H}$ for the water-filling estimator, ice-filling estimator, and random estimator can be respectively written as

### v2

@@ VI-B Performance Evaluation Under Perfect Kernel
In this subsection, we use the perfect kernel, ${\bm{\Sigma}}_{{\mathbf{h}}}$, to trigger the water-filling, ice-filling, and MM algorithms, and then perform the Bayesian regression. The NMSE performance as a function of SNR is evaluated in Fig. 4. The SNR ranges from $-10{\>}{\text{dB}}\sim 15{\>}{\text{dB}}$. Thanks to the careful design of observation matrix, the ideal water-filling, the proposed ice-filling, and the proposed MM algorithms remarkably outperform existing benchmarks. In particular, a 5 dB NMSE gap between ice-filling and MMSE, as well as a 15 dB gap between ice-filling and AMP, are visible when the SNR is 10 dB. Notably, the NMSE performance attained by the ice-filling method tightly aligns with that of the ideal water-filling scheme, given that the ice-filling intrinsically allocates $Q$ quantized units of power to the $Q$-timeslot pilots. Another finding of interest is that the NMSE gap between the MM and ice-filling algorithms is no higher than 1 dB, demonstrating the near-optimality of the proposed MM algorithm in phase-only-controllable design.
@@ VI-C Performance Evaluation Under Imperfect Kernel
In some scenarios where the perfect prior kernel $\bm{\Sigma}_{\bf h}$ cannot be obtained, the statistical kernel and artificial kernels defined in Section IV-E can be utilized to replace the perfect kernel $\bm{\Sigma}_{\bf h}$ in Algorithms 1$\sim$2. We first consider the statistical kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}={\bf{\Sigma}}_{{\bf h}}+\sigma^{2}_{{\mathbf{h}}}{\mathbf{I}}_{M}$. The NMSE as a function of the kernel estimation error $\sigma^{2}_{{\mathbf{h}}}$ is shown in Fig. 7. One can observe that all kernel-dependent methods, including the MMSE, Random ${\mathbf{W}}$, and the proposed estimators, suffer from visible performance loss due to the imperfect kernel input. Fortunately, the proposed methods still hold the superiority among all schemes. In particular, the estimation accuracy of the proposed methods is almost not influenced when $\sigma^{2}_{{\mathbf{h}}}<-15\,{\text{dB}}$, which shows their high robustness to the kernel error. One more interesting finding is that, when $\sigma^{2}_{{\mathbf{h}}}$ increases, the gap between MMSE and our proposed methods, become small. The reason is that, an increasing $\sigma^{2}_{{\mathbf{h}}}$ gradually obscures the structural characteristics of ${\bf{\Sigma}}_{{\bf h}}$, which makes the input kernel ${\bf{\hat{\Sigma}}}_{{\bf h}}$ behaves more like an identity matrix. In this case, the importance of designing the observation matrix $\bf W$ is reduced.
@@
We further evaluate the achievable NMSE performance under the artificial kernels $\bm{\Sigma}_{\rm exp}$ and ${\bm{\Sigma}}_{\rm Bes}$. Their hyper-parameters, $\eta_{1}$ and $\eta_{2}$, are determined by the binary search through several numerical experiments, which are set to $\eta_{1}=0.56$ and $\eta_{2}=0.85$ in the considered scenario. In this way, we obtain the NMSE versus the SNR and that versus the pilot length $Q$ in Fig. 8 and Fig. 9, respectively. From these two figures one find that, for a given kernel, the curves of “ideal water-filling”, “proposed ice-filling”, and “proposed MM” are very close, which indicates that these three schemes have similar robustness to imperfect kernels. This phenomenon proves the feasibility of using experiential kernels.
@@ VII Conclusions
Appendix A Proof of Lemma 1
@@
where $(a)$ comes from the definition of ${\bf\Sigma}_{t-1}$, i.e., ${\bf\Sigma}_{t-1}={{\bf{\Sigma}}_{\bf{h}}}-{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}$. Then, by substituting (56) and (A) into (48), the inverse of ${\bf{K}}_{t}$ can be written as (60).
${\bf{K}}_{t}^{-1}=\left[\begin{array}[]{cc}{\bf{K}}_{t-1}^{-1}+\dfrac{{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}&-\dfrac{{\bf{K}}_{t-1}^{-1}{{\bf{W}}_{t-1}^{H}}{{\bf{\Sigma}}_{\bf{h}}}{\bf{w}}_{t}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}\vspace*{0.3cm}\\ -\dfrac{{\bf{w}}_{t}^{H}{{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}_{t-1}}{\bf{K}}_{t-1}^{-1}}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}&\dfrac{1}{{\bf{w}}_{t}^{H}{\bf\Sigma}_{t-1}{\bf{w}}_{t}+\sigma^{2}}\end{array}\right]$ (60)
@@
It is straightforward to verify that the sum of ${\bf E}$, ${\bf F}$, ${\bf G}$, and ${\bf H}$ is exactly $-{{\bf{\Sigma}}_{t-1}}{\bf w}_{t}{\bf w}_{t}^{H}{{\bf{\Sigma}}_{t-1}}$. By substituting ${\bf E}$, ${\bf F}$, ${\bf G}$, and ${\bf H}$ into (A), we have
@@
Appendix B Proof of Theorem 2
@@
where the inequality (a) holds because $\left(\beta-\frac{\sigma^{2}}{\lambda_{k}}\right)^{+}\geq\beta-\frac{\sigma^{2}}{\lambda_{k}}$. Combining equations (63) and (B), we arrive at
@@
We introduce the following Lemma 2.1 to evaluate the inequality (65).
Lemma 2.1: For the $k$-th eigenvalue $\lambda_{k}^{t}$ obtained at the $t$-th timeslot, if $n_{k}^{t}>0$, we have the inequality
@@
Comparing (65) and (66), it is clear that (65) contradicts Lemma 2.1 given that $n_{k}\geq p_{k}+1>0$. Therefore, $n_{k}$ should be smaller than $p_{k}+1$.
@@
which contradicts Lemma 2.1 because $n_{k^{\prime}}>p_{k^{\prime}}\geq 0$. As a result, $n_{k}$ should be larger than $p_{k}-1$, and thereby we get
@@
Appendix C Proof of Lemma 2
@@
Appendix D Proof of Lemma 3
@@
Appendix E Proof of Theorem 3
@@
Appendix F Proof of Lemma 4
@@
Therefore, by substituting (F) into (F), the MSE $\delta_{\rm rnd}$ is given by
@@
Appendix G Proof of Lemma 5
@@
Appendix H Proof of Lemma 6
@@
which is equal to ${\bm{\Omega}}{{\bf{\hat{\Sigma}}}}_{\bf{h}}$ and obviously the is a function of ${\bf W}{\bf W}^{H}$. Using the same matrix techniques as (H), we can prove that
@@
Appendix I Proof of Lemma 7
Let ${\bf\Sigma_{h}}={\mathbf{U}}{\bf\Lambda}{\mathbf{U}}^{H}$ denote the complete eigenvalue decomposition of the kernel ${{\bf{\Sigma}}_{\bf{h}}}$, wherein ${\mathbf{U}}\in{\mathbb{C}}^{M\times M}$ and ${\mathbf{\Lambda}}=\text{diag}\{\lambda_{1},\cdots,\lambda_{M}\}$ with $\lambda_{m}=0$ for $K<m\leq M$. According to the analyses in Appendices C, D, and F, the term ${\bf W}{\bf W}^{H}$ for the water-filling estimator, ice-filling estimator, and random estimator can be respectively written as